OpenStudy (anonymous):

A room has ana area of 189 square feet. If the room is square how far would an ant walk around to get to the perimeter?

OpenStudy (anonymous):

68

OpenStudy (anonymous):

side length = sqrt(189)= 3sqrt21==> the perimeter =4*3sqrt21 =12sqrt(21)